# Compression fronts from fast radio bursts

### v1

@@ 6 Relaxation to the steady state
A simple argument shows that the timescale to approach the steady state is long if the plasma magnetization $\sigma_{\rm u}$ is large. Consider the vacuum limit of $\rho_{\rm u}\rightarrow 0$ (which corresponds to $\sigma_{\rm u}\rightarrow\infty$). Then, the background field must remain unchanged from its upstream state: in a vacuum, the electromagnetic wave goes through the background field without interaction, so no $E_{\rm bg}$ is generated. Thus, at $\sigma_{\rm u}\rightarrow\infty$ the magnetized plasma must remain static and will not be compressed by the radio wave packet. The steady-state front with large $C$ still provides a consistent solution, but relaxing to it takes an infinite time, so it becomes irrelevant at $\sigma_{\rm u}\rightarrow\infty$.
@@
A similar conclusion is reached by considering energy. Compression of the magnetically dominated plasma requires deposition of energy $\propto\sigma_{\rm u}$, much greater than the energy of plasma particles that interact with the wave. The background field serves as a ballast coupled to the plasma. It increases the effective mass per particle from $m$ to $\sigma_{\rm u}m$, which makes it heavy and impossible to accelerate in the limit of $\sigma_{\rm u}\rightarrow\infty$. An analogy can be drawn with collision of a photon of energy ${\cal E}_{\rm ph}$ with a particle of mass $\sigma m$. Using the frame where the particle is initially at rest, one finds that the particle gains speed $\beta\approx{\cal E}_{\rm ph}/\sigma mc^{2}\ll 1$ if $\sigma\gg{\cal E}_{\rm ph}/mc^{2}$.
It is worth looking in more detail at a similar collision problem formulated for the wave-plasma system.
6.1 Wave-plasma collision
Suppose a radio wave packet of energy ${\cal E}$ passes through an initially static, cold slab of plasma with magnetization $\sigma_{\rm u}$. Some energy $\Delta{\cal E}$ and momentum $\Delta{\cal E}/c$ are lost by the radio wave and received by the slab. As a result, the slab is accelerated to speed $\beta_{\rm D}$ and compressed by the factor $1-\beta_{\rm D}$ (cf. Equation (55)). Compression increases the slab magnetic field from initial $B_{\rm u}$ to $B=B_{\rm u}/(1-\beta_{\rm D})$.
In the simplified collision picture, the slab is treated as a single body characterized by its volume $V$, particle number $N$, inertial mass of the plasma $M=Nm(1+w)$ (increased from $M_{\rm u}=Nm$ by deposited heat), magnetic field $B$, and electric field $E=\beta_{\rm D}B$. Energy conservation gives
$M_{\rm u}c^{2}+V_{\rm u}\frac{B_{\rm u}^{2}}{8\pi}+\Delta{\cal E}=\gamma_{\rm D}Mc^{2}+V\frac{B^{2}+E^{2}}{8\pi}+{\cal E}_{\rm rad},$ (60)
where we included possible radiative losses of the plasma ${\cal E}_{\rm rad}$. The radiated momentum ${\cal P}^{\prime}_{\rm rad}=0$ in the fluid rest frame implies ${\cal P}_{\rm rad}=\beta_{\rm D}{\cal E}_{\rm rad}/c$ in the lab frame. Momentum conservation equation (multiplied by $c$) is
$\Delta{\cal E}=\gamma_{\rm D}\beta_{\rm D}Mc^{2}+V\frac{EB}{4\pi}+\beta_{\rm D}{\cal E}_{\rm rad}.$ (61)
Subtracting the momentum equation from the energy equation one can exclude $\Delta{\cal E}$:
$M_{\rm u}c^{2}+V_{\rm u}\frac{B_{\rm u}^{2}}{8\pi}=\frac{Mc^{2}}{\kappa}+V\frac{(B-E)^{2}}{8\pi}+(1-\beta_{\rm D}){\cal E}_{\rm rad}.$ (62)
We now substitute $E=\beta_{\rm D}B$, $B=B_{\rm u}/(1-\beta_{\rm D})$, $VB=V_{\rm u}B_{\rm u}$, and $M=(1+w)M_{\rm u}$, and obtain
$1=\frac{1+w}{\kappa}-\frac{\beta_{\rm D}\sigma_{\rm u}}{2}+(1-\beta_{\rm D})\epsilon_{\rm rad},$ (63)
where
$\sigma_{\rm u}\equiv\frac{V_{\rm u}B_{\rm u}^{2}}{4\pi M_{\rm u}c^{2}},\qquad\epsilon_{\rm rad}\equiv\frac{{\cal E}_{\rm rad}}{M_{\rm u}c^{2}}.$ (64)
Radiative losses can be described by the effective scattering cross section $\sigma_{\rm sc}$ per electron, with $\sigma_{\rm sc}\sim w^{2}\sigma_{\rm T}$ within a factor of $\sim 2$ (B22). It implies
$\epsilon_{\rm rad}\sim\frac{\sigma_{\rm T}w^{2}\Sigma}{mc^{2}},$ (65)
where $\Sigma$ is the column energy density of the radio wave packet.
Using the identity $1-\beta_{\rm D}=2/(\kappa^{2}+1)$ one can rewrite Equation (63) as a cubic equation for $\kappa$. Its solution simplifies in two limits: $\kappa\gg 1$ ($\beta_{\rm D}\approx 1$) and $\beta_{\rm D}\ll 1$ ($\kappa\approx 1$). Assuming $\sigma_{\rm u}\gg 1$, we find
$\kappa-1\approx\left\{\begin{array}[]{lr}\vspace*{1mm}\displaystyle{\frac{2}{\sigma_{\rm u}}(w+\epsilon_{\rm rad})}&{\rm when}\;\displaystyle{w+\epsilon_{\rm rad}\ll\sigma_{\rm u}}\\ \displaystyle{\frac{w}{\sigma_{\rm u}}+\sqrt{\frac{\epsilon_{\rm rad}}{\sigma_{\rm u}}}}&{\rm when}\;\displaystyle{w+\epsilon_{\rm rad}\gg\sigma_{\rm u}}\end{array}\right.$ (66)
It also gives the proper compression factor $\tilde{C}\approx\kappa$.
Once $\kappa$ is known, we can also find the energy $\Delta{\cal E}$ taken from the radio wave packet. Equation (61) gives
$\frac{\Delta{\cal E}}{M_{\rm u}c^{2}}=\beta_{\rm D}\left[\gamma_{\rm D}(1+w+\sigma_{\rm u}\kappa)+\epsilon_{\rm rad}\right],$ (67)
and using Equation (66) we find
$\frac{\Delta{\cal E}}{M_{\rm u}c^{2}}\approx\left\{\begin{array}[]{lr}\vspace*{1mm}\displaystyle{2(w+\epsilon_{\rm rad})},&\displaystyle{w+\epsilon_{\rm rad}\ll\sigma_{\rm u}}\\ \displaystyle{\frac{w^{2}}{\sigma_{\rm u}}+\frac{3}{2}w\sqrt{\frac{\epsilon_{\rm rad}}{\sigma_{\rm u}}}+\frac{3}{2}\epsilon_{\rm rad}},&\displaystyle{w+\epsilon_{\rm rad}\gg\sigma_{\rm u}}\end{array}\right.$ (68)
The upstream plasma is composed of many slabs. One after another, they fill the propagating packet and exit behind it, and one might conclude that Equation (66) approximately describes $\kappa$ in a steady propagating front. In fact, Equation (66) is very far from the correct MHD steady-state solution given in section 5. When viewed in the upstream rest frame, a non-radiative steady front should have $\kappa=1+w$ regardless of $\sigma_{\rm u}$ (Equation 57). In contrast, the collision model described above yields $\kappa\approx 1+(2w/\sigma_{\rm u})\approx 1$ when $\sigma_{\rm u}\gg w$.
@@ 6.1 Wave-plasma collision
The simplified collision model (Equation 66) works well at the beginning of packet propagation through the plasma; however, it fails to describe the gradual relaxation of the compression front to the correct steady state after many interactions with many plasma slabs. When $\sigma_{\rm u}\gg 1$, this relaxation takes a long time and occurs due to a slow growth of magnetic pressure in the front, as demonstrated below. The grown magnetic pressure drops out from the final steady pattern of the MHD flow because it occurs with a zero gradient of $\sigma/\kappa$ (section 5). This behavior is not captured by the simplified collision model, which treats each plasma slab passing through the packet as a single body and does not consider any gradients. By contrast, in full MHD, the gradients across the front play an essential role: they endow the propagating front with a “memory” of its earlier interactions with the fluid that has been disposed behind the front. The accumulating memory enables the front relaxation to the steady state.
@@
Below we calculate this relaxation by solving the full time-dependent MHD problem described by Equations (50) and (51). We solve it first in the non-radiative regime (neglecting $\dot{\gamma}_{e}$ in Equation (51)) and then in the radiative regime. In the remainder of this section, we assume regular heating ($\tilde{\omega}_{B}>\tilde{\omega}$). A similar relaxation trend holds with stochastic heating ($\tilde{\omega}_{B}<\tilde{\omega}$); however, this regime gives a huge enthalpy $w$ and relaxation takes a very long time, so steady fronts with stochastic heating are not expected in FRBs (section 9).
6.2 Evolution of a non-radiative compression front
@@
[…]
In particular, in a steady state, $\kappa=\kappa_{\star}\approx 1+(3/2)a$. Then, for a packet with $a_{\max}\gg 1$, one finds
[…]
Hereafter, star indicates the steady state.
@@
It grows when the fluid receives momentum from the wave. Recall that the wave interacts with the plasma particles only, and the plasma shares the received momentum with the background field (through gyration). Therefore, the budget for the wave-fluid momentum exchange (per particle crossing the packet) is set by $T^{tz}_{\rm p}$ rather than $T^{tz}_{\rm MHD}$:
@@
where $h=1+w+\sigma$. Let us now specialize to the frame where the upstream fluid is at rest ($\kappa_{\rm u}=1$), and the steady state has $\kappa=\kappa_{\star}=1+w$ and $\sigma=\kappa_{\star}\sigma_{\rm u}=(1+w)\sigma_{\rm u}$. In this frame, the expected timescale to approach the steady state is
$t_{\star}\sim\frac{1+w+\kappa_{\star}\sigma_{\rm u}}{1+w}\,t_{\rm cross}^{\star}=(1+\sigma_{\rm u})t_{\rm cross}^{\star}\quad(\kappa_{\rm u}=1).$ (74)
@@
Figure 3: Top: relaxation of the non-radiative compression front $C(\xi)$ toward the steady-state $C_{\star}(\xi)$ (red curve, Equation (76)) from an initial state with a uniform static plasma, $C=1$ and $\kappa=1$. Black curves show snapshots of the front, with time $t$ indicated next to each curve (in units of packet duration $T$). The wave packet (Equation 5) has $a_{\max}=(4/3)\sqrt{7}$, equivalent to $w_{\max}=4$. The upstream magnetization is $\sigma_{\rm u}=10$. Bottom: evolution of $\kappa$ from the initial $\kappa=1$ toward $\kappa_{\star}$ in the same simulation.
This simple estimate agrees with our numerical solutions for the front evolution. We find the evolution by solving Equations (50) and (51) for $\kappa(t,\xi)$ and $\rho(t,\xi)$ in a given wave packet $a(\xi)$ propagating in a background plasma with a given upstream magnetization $\sigma_{\rm u}$. The fluid enthalpy inside the packet, $w\equiv(U_{\rm p}+P_{\rm p})/\tilde{\rho}c^{2}$, is determined by Equations (9) and (22):
$w(\xi)=\frac{3}{2}\left[\sqrt{a^{2}(\xi)+1}-1\right].$ (75)
@@
In sample models, we use the packet described by Equation (5) with $a_{\max}=(4/3)\sqrt{7}$ which corresponds to $w_{\max}=4$. As an initial state at $t=0$, we take a static plasma, $\kappa(\xi)=1$ and $\rho(\xi)=\rho_{\rm u}$, and follow the evolution of the flow pattern inside the propagating wave packet until it relaxes to the steady-state solution. In the frame where the upstream plasma is at rest ($\kappa_{\rm u}=1$), the steady-state solution is
$\kappa_{\star}=1+w,\quad\rho_{\star}=\frac{\kappa_{\star}^{2}+1}{2}\rho_{\rm u}\qquad(\kappa_{\rm u}=1).$ (76)
@@
The model with $\sigma_{\rm u}=10$ is shown in Figure 3. At the beginning of the simulation we observe that $\kappa(\xi)$ starts to grow from unity and develops a peaked profile — the MHD fluid is accelerated in the first half of the packet and decelerated in the second half, before exiting behind the packet. Over time, $\kappa(\xi)$ approaches $\kappa_{\star}(\xi)$, which is an attractor for the flow pattern. In particular, the maximum of the $\kappa$-profile, $\kappa_{\max}$, eventually approaches the expected $\kappa_{\max}^{\star}=1+w_{\max}=5$. The maximum compression factor $C_{\max}=\rho_{\max}/\rho$ approaches
$C_{\max}^{\star}=\frac{1}{2}\left[(1+w_{\max})^{2}+1\right].$ (77)
@@
The evolution of compression peak $C_{\max}(t)$ provides a simple way to quantify the front relaxation toward the steady-state. The evolution of $C_{\max}$ is shown in Figure 4 for $\sigma_{\rm u}=3$, 10, 20. One can see that the timescale $t_{\star}$ for approaching the steady-state $C_{\max}^{\star}=13$ increases with $\sigma_{\rm u}$ in agreement with Equation (74). The slow evolution occurs through a sequence of quasi-steady states with a nearly uniform mass flux in $\xi$, $F_{\rm m}=\rho(c-v)\approx const$. It implies a correlated growth of $\kappa$ and $C$: the relation $C\approx(\kappa^{2}+1)/2$ is sustained during the evolution (Figure 5).
@@
Figure 5: Relation between $\kappa=\gamma_{\rm D}(1+\beta_{\rm D})$ and $C=\rho/\rho_{\rm u}$ (here measured at the peak inside the compression front) during relaxation toward the final steady state. Black curve shows the simulation result (the model with $\sigma_{\rm u}=20$), and red dashed curve shows the relation $C=(\kappa^{2}+1)/2$, equivalent to $\tilde{\rho}/\kappa=\rho_{\rm u}$ (a uniform mass flux in $\xi$, $F_{\rm m}=\rho(c-v)$). Fluid density in the front closely tracks the steady-state solution $\tilde{\rho}=\kappa\rho_{\rm u}$ defined for an instantaneous profile of $\kappa(\xi)$, which slowly grows during relaxation. The correlated growth of $\kappa$ and $C$ stops when they reach $\kappa_{\max}=5$ and $C_{\max}=13$. Then, the fluxes $F_{\rm m}$ and $F_{-}$ both become exactly uniform across the front, and a true steady state is achieved.
@@
We conclude that at any large but finite $\sigma_{\rm u}$, the wave packet approaches the steady state on a long timescale $t_{\star}\sim\sigma_{\rm u}t_{\rm cross}$ (assuming negligible damping of the radio wave during this time). The conclusion might seem counter-intuitive: one might expect that the front has no memory on timescales longer than $t_{\rm cross}$ since the plasma interacting with the wave packet is continually refreshed on the timescale $t_{\rm cross}$. Nevertheless, the conservation laws require the propagating wave to retain some memory of its previous interactions with the plasma. Given enough time, the packet gradually builds up the strong compression $\tilde{\rho}/\rho_{\rm u}=\kappa\approx w$ — the pattern of the MHD flow in the wave is “groomed” toward the steady state. Remarkably, this gradual evolution leads to a strong magnetic field in the packet, so that the plasma magnetization approaches $\sigma\approx w\sigma_{\rm u}$. Thus, tremendous magnetic pressure develops in the propagating front while it finds a self-consistent steady pattern. At the same time, the final pattern of fluid motion established in the wave packet is the same as it would be in the absence of magnetic pressure, as the background magnetic field drops out from the steady-state solution $\kappa_{\star}(\xi)$.
6.3 Evolution of a radiative compression front

### v2

@@ VI Relaxation to the steady state
A simple argument shows that the approach to the steady state takes the front a long propagation time if the plasma magnetization $\sigma_{\rm u}$ is large. Consider the vacuum limit of $\rho_{\rm u}\rightarrow 0$, which corresponds to $\sigma_{\rm u}\rightarrow\infty$. [^10] Recall that we consider radio waves with frequencies exceeding the plasma Larmor frequency. In the opposite case, they would be MHD waves (where magnetized particles move with $\beta_{\rm D}=E/B$), and $\rho_{\rm u}\rightarrow 0$ would correspond to force-free electrodynamics. Then, the background field must remain unchanged from its upstream state: in a vacuum, the electromagnetic wave goes through the background field without interaction, so no $E_{\rm bg}$ is generated. Thus, at $\sigma_{\rm u}\rightarrow\infty$ the magnetized plasma must remain static and will not be compressed by the radio wave packet. The steady-state front with a large compression factor $C=\rho/\rho_{\rm u}$ still provides a consistent solution, but relaxing to it takes an infinite time, so it becomes irrelevant at $\sigma_{\rm u}\rightarrow\infty$.
@@
The same conclusion is reached by considering energy. Compression of the magnetically dominated plasma requires deposition of energy $\propto\sigma_{\rm u}$, much greater than the energy of plasma particles that interact with the wave. The background field serves as a ballast coupled to the plasma. It increases the effective mass per particle from $m$ to $\sigma_{\rm u}m$, which makes it heavy and impossible to accelerate in the limit of $\sigma_{\rm u}\rightarrow\infty$. An analogy can be drawn with absorption of a photon with energy ${\cal E}_{\rm ph}$ by a particle of mass $\sigma m$. Using the frame where the particle is initially at rest, one finds that the particle gains speed $\beta\approx{\cal E}_{\rm ph}/\sigma mc^{2}\ll 1$ if $\sigma\gg{\cal E}_{\rm ph}/mc^{2}$ (and ${\cal E}_{\rm ph}$ transforms mainly into the particle’s internal energy).
Similarly, one might expect that the interaction of a radio wave packet with the MHD fluid gives it a small bulk velocity,
$\beta_{\rm D}\sim\frac{w}{\sigma_{\rm u}}\ll 1,$ (59)
when the interaction generates fluid enthalpy $w\ll\sigma_{\rm u}$. This would mean that energy deposition by the radio wave heats the fluid without significant bulk acceleration or compression. [^11] A toy problem of wave-plasma collision may be formulated treating the magnetized fluid as a single body carrying energy and momentum. A consistent formulation, however, requires some care: the MHD fluid would need to be confined by walls, which carry stress and energy to be included in the conservation laws. This estimate clearly differs from the steady-front solution (Equation 56), which gives $\kappa\gg 1$ (i.e. $\beta_{\rm D}\approx 1$) at $w\gg 1$ regardless of $\sigma_{\rm u}$.
@@
The estimate $\beta_{\rm D}\sim w/\sigma_{\rm u}$ may be reasonable at the beginning of the packet propagation through the plasma with $\sigma_{\rm u}\gg w$; however, it fails to describe the gradual relaxation of the compression front to the steady state. When $\sigma_{\rm u}\gg 1$, this relaxation takes a long time and involves a slow growth of the compression factor and the magnetic pressure in the front, as demonstrated below. Pressure gradients across the front play an essential role: they endow the propagating front with a “memory” of its earlier interactions with the fluid that has been disposed behind the front. The accumulating memory enables the front relaxation to the steady state.
@@
Below we demonstrate this relaxation by solving the full time-dependent MHD problem described by Equations (46) and (47). We solve it first in the non-radiative regime (neglecting $\dot{{\cal Q}}_{e}^{-}$ in Equation (47)) and then in the radiative regime. In the remainder of this section, we assume regular oscillations of particles in the radio wave. A similar relaxation trend holds with stochastic heating; however, this regime gives a huge enthalpy $w$ and relaxation would take a very long time, so steady fronts with stochastic heating do not form in FRBs (section IX).
VI.1 Evolution of a non-radiative compression front
@@
$t_{\rm cross}=\int_{0}^{T}\frac{d\xi}{1-\beta_{\rm D}}=\int_{0}^{T}\frac{\kappa^{2}+1}{2}\,d\xi.$ (60)
In particular, in a steady state and in the regular oscillation regime, $\kappa^{2}=(1+a^{2})\kappa_{\rm u}^{2}$. Then, for a packet with $a_{\max}\gg 1$, one finds
$t_{\rm cross}^{\star}\approx\int_{0}^{T}[a^{2}(\xi)\,\kappa_{\rm u}^{2}+1]\frac{d\xi}{2}=\left(\frac{3}{8}a_{\max}^{2}\kappa_{\rm u}^{2}+1\right)\frac{T}{2},$ (61)
where we assumed a wave packet $a(\xi)$ of the form given in Equation (7). Hereafter, star indicates a steady state.
@@
It grows when the fluid receives momentum from the radio wave. Recall that the wave directly interacts only with the plasma particles, and the plasma shares the received momentum with the background electromagnetic field (through gyration about $\boldsymbol{B}$). Therefore, the budget for the wave-fluid momentum exchange (per particle crossing the packet) is set by $T^{tz}_{\rm p}$ rather than $T^{tz}_{\rm MHD}$:
@@
where $h=1+w+\sigma$. Let us specialize to the frame where the upstream fluid is at rest ($\kappa_{\rm u}=1$). The steady front has $\kappa=\kappa_{\star}=1+w$ and $\sigma=\kappa_{\star}\sigma_{\rm u}=(1+w)\sigma_{\rm u}$. In this frame,
$t_{\rm cross}^{\star}\approx\frac{3}{16}a_{\max}^{2}T\qquad(a_{\max}\gg 1),$ (65)
and the expected timescale to approach the steady state is
$t_{\star}\sim\frac{1+w+\kappa_{\star}\sigma_{\rm u}}{1+w}\,t_{\rm cross}^{\star}=(1+\sigma_{\rm u})t_{\rm cross}^{\star}.$ (66)
@@
Figure 4: Top: relaxation of the non-radiative compression front $C(\xi)$ toward the steady-state $C_{\star}(\xi)$ (red curve, Equation (67)) from an initial state with a uniform static plasma, $C=1$ and $\kappa=1$. Black curves show snapshots of the front at times $t$ indicated next to each curve in units of packet duration $T$. The wave packet (Equation 7) has $a_{\max}=\sqrt{24}$, which gives $w_{\max}=4$. The upstream magnetization is $\sigma_{\rm u}=10$. Bottom: evolution of $\kappa$ (black) from the initial $\kappa=1$ toward $\kappa_{\star}$ (red) in the same simulation.
@@
This simple estimate agrees with our numerical solution for the front evolution. We find the evolution by solving Equations (46) and (47) for $\kappa(t,\xi)$ and $\rho(t,\xi)$ in a given wave packet $a(\xi)$ propagating in a background plasma with a given upstream magnetization $\sigma_{\rm u}$. The fluid enthalpy inside the packet is given by Equation (45a). In sample models, we use the packet described by Equation (7) with $a_{\max}=\sqrt{24}$, which corresponds to $w_{\max}=4$. As an initial state at $t=0$, we take a static plasma, $\kappa(\xi)=1$ and $\rho(\xi)=\rho_{\rm u}$, and follow the evolution of the flow pattern in the propagating wave packet until it relaxes to a steady-state. In the frame where the upstream plasma is at rest ($\kappa_{\rm u}=1$), the expected steady-state solution is
$\kappa_{\star}=\sqrt{1+a^{2}},\qquad C_{\star}\equiv\frac{\rho_{\star}}{\rho_{\rm u}}=\frac{\kappa_{\star}^{2}+1}{2}=1+\frac{a^{2}}{2}.$ (67)
@@
The sample model with $\sigma_{\rm u}=10$ is shown in Figure 4. At the beginning of the simulation we observe that $\kappa(\xi)$ starts to grow from unity and develops a peaked profile — the MHD fluid is accelerated in the leading part of the packet (at $\xi<\xi_{\star}$) and decelerated in the later part (at $\xi>\xi_{\star}$), before exiting behind the packet. The compression factor in the acceleration part is well described by the steady-state solution:
$\kappa\approx\kappa_{\star}\;\;{\rm and}\;\;C\equiv\frac{\rho}{\rho_{\rm u}}\approx C_{\star}\quad{\rm at~}\xi<\xi_{\star}.$ (68)
The maximum compression factor is reached at $\xi_{\star}$:
$C_{\max}=C(\xi_{\star})\approx C_{\star}(\xi_{\star}).$ (69)
The peak $C_{\max}$ gradually grows with time as its location $\xi_{\star}$ shifts toward the middle of the wave packet.
Figure 5: Relation between $\kappa=\gamma_{\rm D}(1+\beta_{\rm D})$ and $C=\rho/\rho_{\rm u}$ (here measured at the peak inside the compression front) during relaxation toward the final steady state. Black curve shows the simulation result (the model with $\sigma_{\rm u}=10$), and red dashed curve shows the relation $C=(\kappa^{2}+1)/2$ (equivalent to $\tilde{\rho}/\kappa=\rho_{\rm u}$ and uniform mass flux in $\xi$, $F_{\rm m}=\rho(c-v)$). Fluid density in the front closely tracks the steady-state solution $\tilde{\rho}=\kappa\rho_{\rm u}$ defined for an instantaneous profile of $\kappa(\xi)$, which slowly grows during relaxation. The correlated growth of $\kappa$ and $C$ stops when they reach $\kappa_{\max}^{\star}=5$ and $C_{\max}^{\star}=13$. Then, the fluxes of mass and $Q^{-}$ both become exactly uniform across the front, so a true steady state is achieved.
When $\xi_{\star}$ reaches the middle of the wave packet, the entire compression front $\kappa(\xi)$ approaches $\kappa_{\star}(\xi)=\sqrt{1+a^{2}}$, which is an attractor for the flow pattern. In particular, the maximum of the $\kappa$-profile, $\kappa_{\max}$, eventually approaches the expected $\kappa_{\max}^{\star}=1+w_{\max}=5$. The maximum compression factor $C_{\max}=\rho_{\max}/\rho_{\rm u}$ approaches $C_{\max}^{\star}=1+a_{\max}^{2}/2=13$.
We also observe in the simulation that the gradual evolution of the front proceeds through a sequence of quasi-steady states: each snapshot of the front shows a nearly uniform mass flux in $\xi$: $F_{\rm m}=\rho(c-v_{\rm D})\approx const$, so $\rho(1-\beta_{\rm D})\approx\rho_{\rm u}$ and $C\approx(1-\beta_{\rm D})^{-1}$. Thus, the relation $C\approx(\kappa^{2}+1)/2$ is sustained during the evolution, i.e. $\kappa$ and $C$ grow in a correlated way (Figure 5).
@@
Evolution of the compression peak $C_{\max}(t)$ provides a simple way to quantify the front relaxation toward the final, exact steady-state. The evolution of $C_{\max}$ is shown in Figure 6 for $\sigma_{\rm u}=3$, 10, 20. One can see that the timescale $t_{\star}$ for approaching the steady-state $C_{\max}^{\star}=13$ increases with $\sigma_{\rm u}$ in agreement with Equation (66).
@@
We conclude that at any large but finite $\sigma_{\rm u}$, the wave packet approaches the steady state on a long timescale $t_{\star}\sim\sigma_{\rm u}t_{\rm cross}^{\star}$ (assuming negligible damping of the radio wave during this time). This conclusion may seem counter-intuitive: one might expect that the front has no memory on timescales longer than $t_{\rm cross}$ since the plasma interacting with the wave packet is continually refreshed on the timescale $t_{\rm cross}$. Nevertheless, the conservation laws require the propagating front to retain some memory of its earlier structure. Given enough time, the front gradually builds up the strong compression factor $\tilde{\rho}/\rho_{\rm u}=\kappa\approx 1+w$ across the entire front — the pattern of the MHD flow in the wave is “groomed” toward the steady state. Remarkably, this gradual evolution leads to a strong magnetic field in the packet, as the plasma magnetization approaches $\sigma\approx(1+w)\sigma_{\rm u}$. Thus, huge magnetic pressure develops in the propagating front while it finds a self-consistent steady pattern. At the same time, the final pattern of fluid motion in the wave packet is the same as found for radio waves propagating in an unmagnetized plasma. The magnetic pressure drops out from the steady-state solution $\kappa=\kappa_{\star}(\xi)$ because $\sigma/\kappa$ is exactly uniform across the compression front (section V).
The relaxation of the compression front can be viewed as the evolution of a driven MHD wave in the plasma. The FRB drives a smooth compression wave of width $cT$, and it takes some time $t\sim t_{\star}$ for the wave pattern to find its equilibrium shape that propagates together with its driver without further change. Recall that in MHD compression tends to be redistributed across the plasma with Lorentz factor $\tilde{\gamma}_{\rm s}\approx\sqrt{1+\sigma_{\rm eff}}$ (Appendix A). The compression pattern evolves due to the driving effect of the FRB and the communication of MHD information across the pattern, which occurs along characteristics $\cal C$$^{-}$ and $\cal C$$^{+}$. The $\cal C$$^{\pm}$ characteristics propagate in $z$ with speeds $\tilde{\beta}_{\pm}=\pm\tilde{\beta}_{\rm s}$ relative to the fluid, which itself moves with speed $\beta_{\rm D}$ relative to a fixed lab frame. The $\cal C$$^{\pm}$ speeds in the lab frame $\beta_{\pm}$ are determined by the relativistic velocity transformation of $\tilde{\beta}_{\pm}$ from the fluid frame $\tilde{\cal K}$ to the lab frame. Instead of speeds $\beta_{\pm}$, it is convenient to use $\kappa_{\pm}\equiv\gamma_{\pm}(1+\beta_{\pm})$, similar to $\kappa$ defined for the fluid speed $\beta_{\rm D}$ in Equation (10). Then, the transformation between the frames takes the form
[…]
Note that $\tilde{\kappa}_{+}\tilde{\kappa}_{-}=1$, since $\tilde{\beta}_{+}=\tilde{\beta}_{\rm s}=-\tilde{\beta}_{-}$. Magnetically dominated plasma with $\sigma_{\rm eff}\gg 1$ has $\tilde{\gamma}_{\rm s}\approx\sqrt{\sigma_{\rm eff}}$, $\tilde{\kappa}_{+}\approx 2\tilde{\gamma}_{\rm s}$ and $\tilde{\kappa}_{-}\approx(2\tilde{\gamma}_{\rm s})^{-1}$.
The MHD characteristics $\cal C$$^{\pm}$ drift in $\xi$ with rate
$\frac{d\xi_{\pm}}{dt}=1-\beta_{\pm}=\frac{2}{\kappa_{\pm}^{2}+1}.$ (71)
Note that MHD information flows across the compression front in the direction of increasing $\xi$: $d\xi_{\pm}/dt>0$. As one can see in Figure 4, the front evolution is basically the gradual spreading of the relaxed zone $\kappa=\kappa_{\star}$ from the leading edge $\xi=0$ toward larger $\xi$. The steady state is first established near the leading edge of the front where the relaxation time $t_{\star}(\xi)$ is short because $t_{\rm cross}(\xi)$ is short for a small leading part $\xi\ll T$. The middle part of the front takes longer to relax; however, its difference from the steady-state solution, $\kappa\neq\kappa_{\star}$, cannot perturb the already relaxed leading part, since MHD information does not spread to smaller $\xi$.
In a plasma with $\sigma_{\rm eff}\gg 1$, MHD information carried by $\cal C$$^{+}$ crosses the front much slower than $\cal C$$^{-}$:
$\frac{d\xi_{+}}{dt}\approx\frac{2}{\kappa_{+}^{2}}\approx\frac{1}{2\sigma_{\rm eff}\kappa^{2}}\ll\frac{d\xi_{-}}{dt}.$ (72)
Thus, the timescale for complete MHD relaxation is related to $\cal C$$^{+}$ propagation across the front rather than $\cal C$$^{-}$. The drift $d\xi_{+}/dt\approx 2/\kappa_{+}^{2}$ is slower than the plasma drift $d\xi/dt\approx 2/\kappa^{2}$ by the factor of $\tilde{\kappa}_{+}^{2}\approx 4\sigma_{\rm eff}$ (when using a lab frame where the plasma has $\kappa\gg 1$). Therefore, the timescale $t_{+}$ for MHD information to cross the compression front is $\approx 4\sigma_{\rm eff}$ times longer than the timescale for the plasma to cross the front:
$t_{+}\approx 4\sigma_{\rm eff}t_{\rm cross}.$ (73)
When the entire front approaches the steady state, $\kappa\approx\kappa_{\star}=\tilde{\gamma}\kappa_{\rm u}$ implies $\sigma_{\rm eff}\approx\sigma_{\rm u}$ (uniform across the front) and $t_{+}^{\star}\approx 4\sigma_{\rm u}t_{\rm cross}^{\star}$, where star indicates a steady state. This expression is similar to the heuristic estimate for relaxation time $t_{\star}$ given in Equation (66), within the numerical factor of $4$.
VI.2 Evolution of a radiative compression front